# A polynomial bound on the pathwidth of graphs edge-coverable by $k$ shortest paths

### v1

@@ 4.1 Two Paths
As a warm-up, we start describing the structure of graphs edge-coverable by $2$ shortest paths, implying an immediate proof of Theorem˜1.3 when $k=2$.
@@
  • for $i\in\{0,\ell\}$, there exist two internally disjoint paths $Q_{i},R_{i}$ (which do not necessarily have the same length), that are separated from $G-(Q_{i}\cup R_{i})$ by the set $\{x_{i}\}$.
See Figure˜9 for an illustration of a skewer. For each $i\in\{0,\ldots,\ell\}$, the graph $C_{i}:=Q_{i}\cup R_{i}$ is called the piece of $x_{i}$. Note that we allow the paths $Q_{i}$ and $R_{i}$ to have length $1$, in which case the graph $C_{i}$ is just the edge $x_{i}x_{i+1}$ (recall that the graphs we consider are simple), while if $Q_{i}$ and $R_{i}$ have length at least $2$, then $C_{i}$ is an even cycle. Note that every skewer $G$ has a $2$-layering, and thus pathwidth at most $2$, and that moreover, for every $i\in\{0,\ldots,\ell\}$, and every pair of vertices $a,b$ with $a\in V(Q_{i})$ and $b\in V(R_{i})$, there exists a path-decomposition of $G$ of width $2$ in which there exists a bag equals to $\{a,b\}$.
@@ 4.2 Three Paths
This subsection consists in a proof of Theorem˜1.3 when $k=3$, which we restate here for convenience.
@@
Assume without loss of generality that $V(P_{1})\cap V(P_{2})\subseteq V(P_{3})$, and that $G$ is connected. Note that if one of the two paths $P_{1},P_{2}$ is disjoint from $P_{3}$, then $G$ is not connected and we get a contradiction, hence both paths intersect $P_{3}$. For each $i\in\{1,2\}$, we let $u_{i},v_{i}$ denote the two vertices from $V(P_{i})\cap V(P_{3})$ maximizing $d(u_{i},v_{i})$, and such that $d(a_{3},u_{i})\leq d(a_{3},v_{i})$, and we let $B_{i}:=P_{i}[u_{i},b_{i}]$ (see Figure˜10). Then the graph $H:=P_{3}\cup B_{1}\cup B_{2}$ admits a $3$-layering $(V_{0},\ldots,V_{r})$, where $r:=\mathrm{ecc}_{H}(a_{3})$, and for each $i\in\{0,\ldots,r\}$, $V_{i}:=\{u\in V(H):d(a_{3},u)=i\}$; this follows from the observation that the three paths $P_{3},B_{1},B_{2}$ are vertical with respect to $a_{3}$ (note that the hypothesis that $V(P_{3})\subseteq V(P_{1})\cap V(P_{2})$ is used here). In particular, note that this layering satisfies the property that each path $P_{j}$ intersects at most once each layer $V_{i}$. Observe now that $G$ is obtained from $H$ by attaching the two pendant paths $A_{1}:=P_{1}[a_{1},u_{1}]$ and $A_{2}:=P_{2}[a_{2},u_{2}]$ to the vertices $u_{1}$ and $u_{2}$. We claim that it is then not hard to construct a $3$-layering of $G$ obtained from the one of $H$ by adding these two pendant paths to the layering $(V_{i})_{0\leq i\leq r}$. ∎
@@
We now set $H:=P_{1}\cup P_{2}$. By Proposition˜4.3, $H$ is a skewer, and we let $x_{1},\ldots,x_{\ell}$, $R_{0},\ldots,R_{\ell}$, $Q_{0},\ldots,Q_{\ell}$, denote its corresponding components with the notations of Definition˜4.1. Assume first that there exists some $i\in\{0,\ell\}$, such that $a\in R_{i}$ and $b\in Q_{i}$. Then $H$ has a path-decomposition of width $2$ that contains a bag equal to $\{a,b\}$. In particular, as $\{a,b\}$ separates $V(P_{3})$ from $V(H)$ in $G$, we can insert after the bag $\{a,b\}$ of this path-decomposition some path-decomposition of width $1$ of $P_{3}$ to which we add in each bag the set $\{a,b\}$. This gives a path-decomposition of width at most $3$ of $G$.
@@
To conclude the proof of Theorem˜1.3, we will prove that $\mathrm{pw}(G)\leq 3$ in each of these cases in the following separate lemmas. This is the most technical part of the proof, especially for cases (1) and (2). The next four lemmas respectively deal with each of the four cases. We start with (1).
@@
All previous paragraphs together with Remark˜4.8 thus imply that we can write $G=G_{1}\cup G_{2}$, where $G_{1}:=A_{1}\cup A_{2}\cup A_{3}$ and $G_{2}:=B_{1}\cup B_{2}\cup B_{3}\cup C_{3}$, and $\{u_{0},u_{1}\}$ separates $G_{1}$ from $G_{2}$. We now show that $G_{1}$ (resp. $G_{2}$) admits a path-decomposition of width at most $3$ whose rightmost (resp. leftmost) bag contains $\{u_{0},u_{1}\}$. This will immediately imply that $\mathrm{pw}(G)\leq 3$.
We first construct such a path-decomposition $\mathcal{P}_{2}$ for $G_{2}$. Observe that $B_{1},B_{2}$ are both vertical with respect to $u_{0}$. Moreover, recall that $C_{3}-u_{3}$ is disjoint from $V(B_{1})\cup V(B_{2})$, hence, as $B_{3}$ is vertical with respect to $u_{0}$, it implies that $B_{3}\cdot C_{3}$ is also vertical with respect to $u_{0}$. It thus implies that the partition $(V_{i})_{i\geq 0}$ of $V(G_{2})$ defined by setting for each $i\in\{0,\ldots,r\}$ (where $r:=\mathrm{ecc}_{G_{2}}(u_{0})$), $V_{i}:=\{x\in V(G_{2}):d(u_{0},x)=i\}$ is a $3$-layering of $G_{2}$. Moreover, note that if we set $s:=d(u_{0},u_{1})$, then the two subsequences $\mathcal{S}_{1}:=(V_{0},\ldots,V_{s})$ and $\mathcal{S}_{2}:=(V_{s},\ldots,V_{r})$ of $(V_{i})_{0\leq i\leq r}$ are respectively $2$- and $3$-layerings of the subgraphs $G^{\prime}_{2},G^{\prime\prime}_{2}$ of $G_{2}$ induced respectively by $V_{0}\cup\cdots\cup V_{s}$ and by $V_{s}\cup\cdots\cup V_{r}$ (see Figure˜13). In particular, two applications of Lemma˜2.3 give path-decompositions $\mathcal{P}^{\prime}_{2},\mathcal{P}^{\prime\prime}_{2}$ respectively of $G^{\prime}_{2}$ and $G^{\prime\prime}_{2}$ such that:
@@
To obtain a path-decomposition $\mathcal{P}_{2}$ of $G_{2}$ with the desired properties, we can now simply add the vertex $u_{1}$ in every bag of $\mathcal{P}^{\prime}_{2}$, and glue the obtained path-decomposition with $\mathcal{P}^{\prime\prime}_{2}$.
@@
It remains to construct a path-decomposition $\mathcal{P}_{1}$ of width $3$ of $G_{1}$ whose rightmost bag contains $\{u_{0},u_{1}\}$. By Remark˜4.8, $\{u_{0}\}$ separates the vertices of $A_{1}$ from the other vertices of $G_{1}$. Moreover, by Proposition˜4.3, $A_{2}\cup A_{3}$ is a skewer. In particular, it admits a path-decomposition $\mathcal{P}^{\prime}_{1}$ of width at most $2$ whose rightmost bag is $\{u_{1}\}$. As $G_{1}$ is obtained from $A_{2}\cup A_{3}$ after attaching some pendant path $A_{1}$ on $u_{0}$, the existence of $\mathcal{P}_{1}$ easily follows. ∎
@@
If $P_{3}$ and $A_{2}$ are disjoint, then $\mathrm{pw}(G)\leq 3$.
@@
Now, observe that Claims 4.11 and 4.12, together with the definition of $v_{1},v_{2}$ imply that we may assume that the set $\{v_{1},v_{2}\}$ separates in $G$ the subgraph $G_{1}:=A_{1}\cdot C_{1}\cup A_{2}\cdot C_{2}$ from the subgraph $G_{2}:=D_{1}\cup E_{1}\cup D_{2}\cup E_{2}\cup P_{3}$ (see Figure˜14). By Proposition˜4.3, $G_{1}$ is a skewer, and admits a path-decomposition $\mathcal{P}^{\prime}_{1}$ of width at most $2$ such that $v_{1}$ is in the rightmost bag of $\mathcal{P}^{\prime}_{1}$. By adding $v_{2}$ to every bag of $\mathcal{P}_{1}^{\prime}$, we obtain a path-decomposition $\mathcal{P}_{1}$ of $G_{1}$ of width at most $3$ such that $v_{1},v_{2}$ is the rightmost bag of $\mathcal{P}_{1}$. It thus remains to show that there exists a path-decomposition of width $3$ of $G_{2}$ whose leftmost bag is $\{v_{1},v_{2}\}$.
@@
Note that by Claims 4.11 and 4.12, the only possible common vertices between $P_{1}$ and $P_{2}$ in $G_{2}$ are in $V(D_{1})\cap V(D_{2})$, and possibly, the vertices $v_{1},v_{2}$ in the special case when $v_{1}=v_{2}$. In particular, by definition of $w_{1},w_{2}$, the set $\{w_{1},w_{2}\}$ separates in $G_{2}$ the vertices of the graph $G_{3}:=D_{1}\cup D_{2}$ from the remaining vertices. By Proposition˜4.3, and as $u_{0}\in V(P_{1})\cap V(P_{2})$, $G_{3}$ admits a path-decomposition of width at most $2$ whose leftmost bag is $\{w_{1},w_{2}\}$. To conclude our proof, it thus only remains to show in each case that there exists a path-decomposition of $G_{4}:=P_{3}\cup E_{1}\cup E_{2}$ of width at most $3$, whose leftmost bag is $\{v_{1},v_{2}\}$, and whose rightmost bag is $\{w_{1},w_{2}\}$. We claim that it is then not hard to construct separately in each case such a decomposition. ∎
@@ 5 Coverings by Isometric Subtrees
A tree-decomposition of a graph $G$ is a pair $(T,\mathcal{V})$ where $T$ is a tree and $\mathcal{V}=(V_{t})_{t\in V(T)}$ is a family of subsets $V_{t}$ of $V(G)$ such that:
@@
The subsets $V_{t}$ are called the bags of the tree-decomposition $(T,\mathcal{V})$, and the tree $T$ is called the decomposition tree. The width of a tree-decomposition is the maximum size of a bag minus $1$. The treewidth of a graph $G$, denoted $\mathrm{tw}(G)$ is defined as the minimum possible width of a tree-decomposition of $G$. Note that path-decompositions of $G$ correspond exactly to the tree-decompositions of $G$ whose decomposition tree is a path. In particular, we have in general $\mathrm{tw}(G)\leq\mathrm{pw}(G)$.
@@
We now prove our main result in this section, which we restate. See 1.4
@@
  • there exists a graph isomorphism $\iota$ from $T$ to $T^{\prime}$ such that for every $v\in V(T_{1})\cap V(T_{2})$, we have $\iota(v)=v$.
The proof of Theorem˜1.4 will be an immediate consequence of the following two Claims.
@@
We let $H$ be a graph with a mirror-decomposition and $T,T^{\prime},\iota$ be as in the definition of mirror-decomposition. We prove by induction on $|V(T)|=|V(T^{\prime})|\geq 1$ that there exists a tree-decomposition of $H$ of width at most $2$, and such for each $v\in V(T)$, there exists a bag that contains $\{v,\iota(v)\}$. If $|V(T)|\leq 2$, then $|V(H)|\leq 3$ (unless $V(T)$ and $V(T^{\prime})$ are disjoint, in which case we also conclude immediately), and one can conclude by considering the tree-decomposition with a single bag equal to $V(H)$. Assume now that $|V(T)|=|V(T^{\prime})|\geq 3$. In particular, there exists some node $u\in V(T)$ such that TuT-u is not connected. Then, $\{u,\iota(u)\}$ forms a separator of size $2$ in $H$ (see Figure˜18).

### v2

@@ 4.1 Two Paths
As a warm-up, we start describing the structure of graphs edge-coverable by $2$ shortest paths, implying an immediate proof of Theorem˜1.4 when $k=2$.
@@
  • there exist four internally disjoint paths $Q_{0},R_{0},Q_{\ell},R_{\ell}$ (which do not necessarily have the same length), such that $Q_{0},R_{0}$ (resp. $Q_{\ell},R_{\ell}$) are separated from $G-(Q_{0}\cup R_{0})$ (resp. $G-(Q_{\ell}\cup R_{\ell})$) by the set $\{x_{1}\}$ (resp. $\{x_{\ell}\}$).
See Figure˜9 for an illustration of a skewer. For each $i\in\{0,\ldots,\ell\}$, the graph $C_{i}:=Q_{i}\cup R_{i}$ is called the piece of $x_{i}$. Note that we allow the paths $Q_{i}$ and $R_{i}$ to have length $1$, in which case the graph $C_{i}$ is just the edge $x_{i}x_{i+1}$ (recall that the graphs we consider are simple), while if $Q_{i}$ and $R_{i}$ have length at least $2$, then $C_{i}$ is an even cycle. Note that every skewer $G$ has a $2$-layering, and thus pathwidth at most $2$, and that moreover, for every $i\in\{0,\ldots,\ell\}$, and every pair of vertices $a,b$ with $a\in V(Q_{i})$ and $b\in V(R_{i})$, there exists a path decomposition of $G$ of width $2$ in which there exists a bag equals to $\{a,b\}$.
@@ 4.2 Three Paths
This subsection consists in a proof of Theorem˜1.4 when $k=3$, which we restate here for convenience.
@@
Assume without loss of generality that $V(P_{1})\cap V(P_{2})\subseteq V(P_{3})$, and that $G$ is connected. Note that if one of the two paths $P_{1},P_{2}$ is disjoint from $P_{3}$, then $G$ is not connected, and we get a contradiction, hence both paths intersect $P_{3}$. For each $i\in\{1,2\}$, we let $u_{i},v_{i}$ denote the two vertices from $V(P_{i})\cap V(P_{3})$ maximizing $d(u_{i},v_{i})$, and such that $d(a_{3},u_{i})\leq d(a_{3},v_{i})$, and we let $B_{i}:=P_{i}[u_{i},b_{i}]$ (see Figure˜10). Then the graph $H:=P_{3}\cup B_{1}\cup B_{2}$ admits a $3$-layering $(V_{0},\ldots,V_{r})$, where $r:=\mathrm{ecc}_{H}(a_{3})$, and for each $i\in\{0,\ldots,r\}$, $V_{i}:=\{u\in V(H):d(a_{3},u)=i\}$; this follows from the observation that the three paths $P_{3},B_{1},B_{2}$ are vertical with respect to $a_{3}$ (note that the hypothesis that $V(P_{3})\subseteq V(P_{1})\cap V(P_{2})$ is used here). In particular, note that this layering satisfies the property that each path $P_{j}$ intersects at most once each layer $V_{i}$. Observe now that $G$ is obtained from $H$ by attaching the two pendant paths $A_{1}:=P_{1}[a_{1},u_{1}]$ and $A_{2}:=P_{2}[a_{2},u_{2}]$ to the vertices $u_{1}$ and $u_{2}$. We claim that it is then not hard to construct a $3$-layering of $G$ obtained from the one of $H$ by adding these two pendant paths to the layering $(V_{i})_{0\leq i\leq r}$. ∎
@@
We now set $H:=P_{1}\cup P_{2}$. By Proposition˜4.3, $H$ is a skewer, and we let $x_{1},\ldots,x_{\ell}$, $R_{0},\ldots,R_{\ell}$, $Q_{0},\ldots,Q_{\ell}$, denote its corresponding components with the notations of Definition˜4.1. Assume first that there exists some $i\in\{0,\ell\}$, such that $a\in R_{i}$ and $b\in Q_{i}$. Then $H$ has a path decomposition of width $2$ that contains a bag equal to $\{a,b\}$. In particular, as $\{a,b\}$ separates $V(P_{3})$ from $V(H)$ in $G$, we can insert after the bag $\{a,b\}$ of this path decomposition some path decomposition of width $1$ of $P_{3}$ to which we add in each bag the set $\{a,b\}$. This gives a path decomposition of width at most $3$ of $G$.
@@
To conclude the proof of Theorem˜1.4, we will prove that $\mathrm{pw}(G)\leq 3$ in each of these cases in the following separate lemmas. This is the most technical part of the proof, especially for cases (1) and (2). The next four lemmas respectively deal with each of the four cases. We start with (1).
@@
All previous paragraphs together with Remark˜4.8 thus imply that we can write $G=G_{1}\cup G_{2}$, where $G_{1}:=A_{1}\cup A_{2}\cup A_{3}$ and $G_{2}:=B_{1}\cup B_{2}\cup B_{3}\cup C_{3}$, and $\{u_{0},u_{1}\}$ separates $G_{1}$ from $G_{2}$. We now show that $G_{1}$ (resp. $G_{2}$) admits a path decomposition of width at most $3$ whose rightmost (resp. leftmost) bag contains $\{u_{0},u_{1}\}$. This will immediately imply that $\mathrm{pw}(G)\leq 3$.
We first construct such a path decomposition $\mathcal{P}_{2}$ for $G_{2}$. Observe that $B_{1},B_{2}$ are both vertical with respect to $u_{0}$. Moreover, recall that $C_{3}-u_{3}$ is disjoint from $V(B_{1})\cup V(B_{2})$, hence, as $B_{3}$ is vertical with respect to $u_{0}$, it implies that $B_{3}\cdot C_{3}$ is also vertical with respect to $u_{0}$. It thus implies that the partition $(V_{i})_{i\geq 0}$ of $V(G_{2})$ defined by setting for each $i\in\{0,\ldots,r\}$ (where $r:=\mathrm{ecc}_{G_{2}}(u_{0})$), $V_{i}:=\{x\in V(G_{2}):d(u_{0},x)=i\}$ is a $3$-layering of $G_{2}$. Moreover, note that if we set $s:=d(u_{0},u_{1})$, then the two subsequences $\mathcal{S}_{1}:=(V_{0},\ldots,V_{s})$ and $\mathcal{S}_{2}:=(V_{s},\ldots,V_{r})$ of $(V_{i})_{0\leq i\leq r}$ are respectively $2$- and $3$-layerings of the subgraphs $G^{\prime}_{2},G^{\prime\prime}_{2}$ of $G_{2}$ induced respectively by $V_{0}\cup\cdots\cup V_{s}$ and by $V_{s}\cup\cdots\cup V_{r}$ (see Figure˜13). In particular, two applications of Lemma˜2.3 give path decompositions $\mathcal{P}^{\prime}_{2},\mathcal{P}^{\prime\prime}_{2}$ respectively of $G^{\prime}_{2}$ and $G^{\prime\prime}_{2}$ such that:
@@
To obtain a path decomposition $\mathcal{P}_{2}$ of $G_{2}$ with the desired properties, we can now simply add the vertex $u_{1}$ in every bag of $\mathcal{P}^{\prime}_{2}$, and glue the obtained path decomposition with $\mathcal{P}^{\prime\prime}_{2}$.
@@
It remains to construct a path decomposition $\mathcal{P}_{1}$ of width $3$ of $G_{1}$ whose rightmost bag contains $\{u_{0},u_{1}\}$. By Remark˜4.8, $\{u_{0}\}$ separates the vertices of $A_{1}$ from the other vertices of $G_{1}$. Moreover, by Proposition˜4.3, $A_{2}\cup A_{3}$ is a skewer. In particular, it admits a path decomposition $\mathcal{P}^{\prime}_{1}$ of width at most $2$ whose rightmost bag is $\{u_{1}\}$. As $G_{1}$ is obtained from $A_{2}\cup A_{3}$ after attaching some pendant path $A_{1}$ on $u_{0}$, the existence of $\mathcal{P}_{1}$ easily follows. ∎
@@
b If $P_{3}$ and $A_{2}$ are disjoint, then $\mathrm{pw}(G)\leq 3$.
@@
Now, observe that Claims 4.11 and 4.12, together with the definition of $v_{1},v_{2}$ imply that we may assume that the set $\{v_{1},v_{2}\}$ separates in $G$ the subgraph $G_{1}:=A_{1}\cdot C_{1}\cup A_{2}\cdot C_{2}$ from the subgraph $G_{2}:=D_{1}\cup E_{1}\cup D_{2}\cup E_{2}\cup P_{3}$ (see Figure˜14). By Proposition˜4.3, $G_{1}$ is a skewer, and admits a path decomposition $\mathcal{P}^{\prime}_{1}$ of width at most $2$ such that $v_{1}$ is in the rightmost bag of $\mathcal{P}^{\prime}_{1}$. By adding $v_{2}$ to every bag of $\mathcal{P}_{1}^{\prime}$, we obtain a path decomposition $\mathcal{P}_{1}$ of $G_{1}$ of width at most $3$ such that $v_{1},v_{2}$ is the rightmost bag of $\mathcal{P}_{1}$. It thus remains to show that there exists a path decomposition of width $3$ of $G_{2}$ whose leftmost bag is $\{v_{1},v_{2}\}$.
@@
Note that by Claims 4.11 and 4.12, the only possible common vertices between $P_{1}$ and $P_{2}$ in $G_{2}$ are in $V(D_{1})\cap V(D_{2})$, and possibly, the vertices $v_{1},v_{2}$ in the special case when $v_{1}=v_{2}$. In particular, by definition of $w_{1},w_{2}$, the set $\{w_{1},w_{2}\}$ separates in $G_{2}$ the vertices of the graph $G_{3}:=D_{1}\cup D_{2}$ from the remaining vertices. By Proposition˜4.3, and as $u_{0}\in V(P_{1})\cap V(P_{2})$, $G_{3}$ admits a path decomposition of width at most $2$ whose leftmost bag is $\{w_{1},w_{2}\}$. To conclude our proof, it thus only remains to show in each case that there exists a path decomposition of $G_{4}:=P_{3}\cup E_{1}\cup E_{2}$ of width at most $3$, whose leftmost bag is $\{v_{1},v_{2}\}$, and whose rightmost bag is $\{w_{1},w_{2}\}$. We claim that it is then not hard to construct separately in each case such a decomposition. ∎
@@ 5 Coverings by Isometric Subtrees
A tree decomposition of a graph $G$ is a pair $(T,\mathcal{V})$ where $T$ is a tree and $\mathcal{V}=(V_{t})_{t\in V(T)}$ is a family of subsets $V_{t}$ of $V(G)$ such that:
@@
The subsets $V_{t}$ are called the bags of the tree decomposition $(T,\mathcal{V})$, and the tree $T$ is called the decomposition tree. The width of a tree decomposition is the maximum size of a bag minus $1$. The treewidth of a graph $G$, denoted $\mathrm{tw}(G)$ is defined as the minimum possible width of a tree decomposition of $G$. Note that path decompositions of $G$ correspond exactly to the tree decompositions of $G$ whose decomposition tree is a path. In particular, we have in general $\mathrm{tw}(G)\leq\mathrm{pw}(G)$.
@@
We now prove our main result in this section, which we restate. See 1.5
@@
  • there exists a graph isomorphism $\iota$ from $T$ to $T^{\prime}$ such that for every $v\in V(T)\cap V(T^{\prime})$, we have $\iota(v)=v$.
The proof of Theorem˜1.5 will be an immediate consequence of the following two Claims.
@@
We let $H$ be a graph with a mirror-decomposition and $T,T^{\prime},\iota$ be as in the definition of mirror-decomposition. We prove by induction on $|V(T)|=|V(T^{\prime})|\geq 1$ that there exists a tree decomposition of $H$ of width at most $2$, and such for each $v\in V(T)$, there exists a bag that contains $\{v,\iota(v)\}$. If $|V(T)|\leq 2$, then $|V(H)|\leq 3$ (unless $V(T)$ and $V(T^{\prime})$ are disjoint, in which case we also conclude immediately), and one can conclude by considering the tree decomposition with a single bag equal to $V(H)$. Assume now that $|V(T)|=|V(T^{\prime})|\geq 3$. In particular, there exists some node $u\in V(T)$ such that TuT-u is not connected. Then, $\{u,\iota(u)\}$ forms a separator of size $2$ in $H$ (see Figure˜18).